OpenStudy (anonymous):

How far back in the taxonomic categories do you have to go to find a group that has living relatives of humans? A. Genus (Homo) B. Family (Hominid) C. Order (primates) D. Class (mammals)

OpenStudy (caominhim):

well lets see F= ma, no that doesn't help V= x/t , nope, oh maybe a derivative A = v/t, no all i'm getting are numbers lets try energy K = 1/2 mv^2, what's the mass of your primate?